### v1

# Context versus Prior Knowledge in Language Models

Abstract
To answer a question, language models often need to integrate prior knowledge learned during pretraining and new information presented in context. We hypothesize that models perform this integration in a predictable way across different questions and contexts: models will rely more on prior knowledge for questions about entities (e.g., persons, places, etc.) that they are more familiar with due to higher exposure in the training corpus, and be more easily persuaded by some contexts than others. To formalize this problem, we propose two mutual information-based metrics to measure a model's dependency on a context and on its prior about an entity: first, the persuasion score of a given context represents how much a model depends on the context in its decision, and second, the susceptibility score of a given entity represents how much the model can be swayed away from its original answer distribution about an entity. Following well-established measurement modeling methods, we empirically test for the validity and reliability of these metrics. Finally, we explore and find a relationship between the scores and the model's expected familiarity with an entity, and provide two use cases to illustrate their benefits.
1 Introduction
Language models have displayed remarkable abilities to answer factual queries about entities, suggesting that they encode knowledge about these entities learned during pretraining (Petroni et al., 2019; Brown et al., 2020; Roberts et al., 2020; Geva et al., 2021). For prompts that extend a question with additional information or context, the model can draw on both its prior knowledge and the additional context in order to answer the query (Kwiatkowski et al., 2019; Joshi et al., 2017; Berant et al., 2013; Kasai et al., 2023). While previous research has investigated how often a model will rely on prior knowledge over conflicting contextual information in answering questions (Longpre et al., 2021), we hypothesize that models will not behave identically for all contexts and entities. For example, if a language model is prompted with Harry hugged Voldemort. How friendly are Harry Potter and Lord Voldemort?, we might expect the prior knowledge learned from training data describing the rivalry between these two characters to significantly influence the model's answer. However, if the model lacks a strong prior on, say, Susie and Alia, then we might expect its answer to be primarily context-driven when prompted with Susie hugged Alia. How friendly are Susie and Alia?.
We formalize this problem through the lens of evaluating the change in a model's answer distribution for different contexts and entities. We present two mutual information-based metrics that allow us to explore differences in the effect of specific contexts on model behavior for different entities. The persuasion score of a given context measures how much a model's answer distribution is affected by the context when prompted with a particular query about a given entity. The susceptibility score of a given entity measures how much the model's answer distribution can be swayed for a particular query about that entity, marginalized over all contexts. Given their basis on mutual information, these metrics are both faithful and comprehensive reflections of model behavior. We offer empirical evidence of the validity and reliability of these metrics by comparing them against similar metrics and showing their robustness (i.e., low variance) to paraphrases and different samples.
Using these metrics and a synthetic dataset111https://github.com/kdu4108/measureLM covering 122 topic areas extracted from the YAGO knowledge graph (Suchanek et al., 2007), we study the effect of a variety of factors on the behavior of pretrained models (ranging from 70m to 12b parameters) across contexts and entities. We find evidence that relevant contexts are consistently more persuasive than irrelevant ones, and assertive contexts are more persuasive than less assertive ones for yes/no questions. In a deep dive for one model, we provide evidence that entities appearing frequently in the training data, and thus that we expect the model to be familiar with, have lower susceptibility scores. Since calculating more complex statistics (i.e., relation-specific entity co-occurrence) from noisy training data can be challenging, we further support our findings by obtaining the same result with a relation-dependent measure of expected familiarity based on the entity's degree in the YAGO knowledge graph.
We further conduct case studies to demonstrate how these metrics could be useful in an applied setting. In a study on friend–enemy stance measurement, we find evidence that enemy duos are less susceptible than friend duos. Applying our metrics to gender bias, we find evidence for a difference in susceptibility between stereotypically masculine and feminine names with respect to gender-biased contexts. Through this, we demonstrate how our proposed metrics can be used to better analyze the effects of context and prior knowledge, with the potential to apply this knowledge to obtain greater control over model behavior.
2 Context and Prior Knowledge
Much prior work has noted that language models develop biases about entities during training and investigated the tension between this entity bias and additional information provided in the context.
2.1 Entity Bias
Loosely defined across studies as model bias from spurious correlations between entity mentions and a target characteristic in the training data, entity bias has been mainly examined in relation extraction (RE), where a model extracts relationships between entities from text (Zelenko et al., 2002; Zhou et al., 2005). Different studies on RE have attempted to either mitigate (Zhang et al., 2017, 2018; Peng et al., 2020; Wang et al., 2022) or leverage (Yamada et al., 2020; Zhou and Chen, 2022) entity bias for improved performance on the task. Wang et al. (2022) note that entities can carry both useful semantic information about their roles (e.g., whether the entity is a person or a place) and spurious information which can bias the model toward relations not mentioned in the target sentence (e.g., the model may link Switzerland with the relation countries_of_residence at inference-time for a sentence that does not mention this relation, due to their frequent co-occurrence in the training data). Additional studies on machine reading comprehension (MRC) tasks, e.g., SQuAD (Rajpurkar et al., 2016), TriviaQA (Joshi et al., 2017), and NaturalQuestions (Kwiatkowski et al., 2019), have found substituting different entity names can result in meaningful changes in model predictions and overall evaluation performance (Yan et al., 2022).
2.2 Context and Entity Bias
The existence of entity bias naturally raises the question of how it interacts with context to shape a model's response. Longpre et al. (2021) approach this by inducing and exploring knowledge conflicts, where a context preceding a query proposes information that conflicts with a model's prior knowledge about that query. They measure the model's reliance on pretrained entity bias and context by computing the memorization ratio: the proportion of knowledge-conflict examples for which the model maintains its answer from prior knowledge. Pezeshkpour (2023) measures a model's prior knowledge of a fact by comparing the entropy of a queried model's answer distribution before and after stating the fact in context. Several studies further propose interventions to reduce entity bias and favor in-context information. These approaches include prompting (Zhou et al., 2023; Onoe et al., 2023), modifying training data (Wang et al., 2023), fine-tuning (Li et al., 2023), and neuron-level interventions during inference (Yu et al., 2023).
However, the metrics used to quantify model reliance on context and entity bias in these works (excepting (Pezeshkpour, 2023)—memorization ratio and empirical task performance on existing RE or MRC tasks—are limited in several ways. First, these works do not find entity-specific or context-specific metrics for how strong the entity bias or context persuasiveness is. Instead, these metrics produce only a single number which summarizes the model's overall reliance on entity bias. Second, these metrics can be unreliable with high variance since they sample from the model's output distribution. Third, their setups are limited to adversarial cases in which a context disagrees with the entity bias. We may well wish to measure the interplay of context and entity bias in other cases, such as when context reinforces entity bias or when they do not clearly disagree. Therefore, we seek to define a metric that measures how much the model depends on a given context or entity with rigorous, theoretically grounded interpretations.
3 Our Formalization
We now formalize the problem setting in which we define metrics for the susceptibility of an entity and the persuasiveness of a context for a given model. Let be an alphabet. Consider a language model over , i.e., is a distribution over . Furthermore, we assume that was estimated from a corpus . Let that corresponds to string representations of entities. Let be a set of slotted query templates of type that fill the slot with the argument to each . We can think of a query template as slotting an entity into a query, e.g., slotting the entity Slovenia into the query template The capital of is produces the string The capital of Slovenia is.222In practice, an entity can have different verbalizations, e.g., Sherlock and Sherlock Holmes. Our formalization does not group different verbalizations into the same entity, which might be useful to explore in future work.
Now, consider three random variables. First, let be a -valued random variable that stands for a context. Second, let be a -valued random variable whose values are answers. Let be a -valued random variable. The pushforward , then, is a -valued random variable over slottings of entities according to query . The random variables , and are jointly distributed according to the following probability
where is string concatenation. To formalize persuasion and susceptibility, we make use of the joint distribution heavily in the subsequent subsections.
3.1 Persuasion Score
Formally, for each context prepended to a query template to an entity , we wish to assign a persuasion score to represent how successful that context is at altering a model's answer distribution. This score depends on the specific queried entity , because contexts themselves are often entity-dependent. Intuitively, a context's persuasion score should measure how much the log probability distribution of possible answers changes, averaged across all possible answers. More precisely, we define our persuasion score as the half-pointwise mutual information (half-PMI) between the context and the answer random variable , conditioned on the fixed query about an entity:
where we define the following marginal distribution
The persuasion score of a context can then be interpreted as the degree (in nats) to which the context was able to change the model's answer distribution when prepended to a query. When the persuasion score is at its lower bound of 0 nats, it indicates the context is completely unpersuasive, i.e., it did not change the model's answer distribution at all. Contexts with higher persuasion scores are more persuasive, i.e., they change the answer distribution more, which is consistent when viewed through the lens of KL-divergence.333We prove further equivalencies of half-PMI and other concepts in information theory in App. A. Due to its information-theoretic foundation, faithfulness to the actual model distribution, and interpretable units, when framed in existing measurement modeling methods, we argue this metric exhibits construct validity via face validity, i.e., the persuasion score is a plausible and meaningful measure for the persuasiveness of a context (Loevinger, 1957; Messick, 1987; Jackman, 2008; Hand, 2004; Quinn et al., 2010; Jacobs and Wallach, 2021).
3.2 Susceptibility Score
For a given query template applied to an entity , we further wish to assign a susceptibility score to which represents how easy it is to change a model's answer distribution. Intuitively, the susceptibility score should measure how much the log probability of an answer, when prompted with just the query, differs from the log probability of that answer when prompted with additional context, averaged across all possible contexts and answers. More precisely, we define the susceptibility score as the mutual information between the context and answer random variables, conditioned on a fixed query about an entity:
where is the marginal distribution over contexts. Equivalent to , the susceptibility score represents the reduction in answer distribution uncertainty (in nats) when a query is preceded by a context, and in a well-grounded statistical sense represents the dependence of answer on context , conditioned on the query–entity pair . A high susceptibility score means the model is highly influenced by context for the query about that entity, with its upper bound of indicating that context fully determines the answer. A low score signals the model's response is robust to context, with its lower bound of indicating no influence of context on the answer distribution. This information-theoretic metric enables us to answer the question ``how much does the model's answer depend on its entity bias?'' with an information-theoretically grounded, interpretable scale based on the model's full behavior. These qualities support our claim that susceptibility scores are meaningful and plausible and, thus, under measurement modeling frameworks, exhibit construct validity and face validity.
3.3 Entity-Independent Persuasion Score
Persuasion scores can be further marginalized over entities. Analogous to our definition of the susceptibility score, we define the entity-independent persuasion score of a context as how much the log probability distribution of possible answers changes, averaged across all possible entities and answers. We describe this further in App. B.
4 Experiments
We now provide empirical evidence to further validate our metrics, characterize model behavior using the susceptibility and persuasion scores, and investigate reasons behind differences in susceptibility scores for different entities.
4.1 Setup
For each of the 122 different relations from the YAGO knowledge graph, we collect 100 entities44450 are real entities (e.g., Adele) sampled from YAGO, and 50 are fake entities (e.g., Udo König) of the same entity class (e.g., Person) generated with GPT-4 (OpenAI, 2023). and construct 600 random contexts from relation-specific context templates such that each entity is mentioned in 6 contexts. The 600 contexts are evenly distributed between assertive,555E.g., Definitely, the capital of {entity} is {answer}. base,666E.g., The capital of {entity} is {answer}. and negation777E.g., The capital of {entity} is not {answer}. context types. We construct four query forms of each relation: two closed questions (i.e., yes–no) and two open questions.
Using these samples, we compute persuasion scores for each context according to Eq. 2 and susceptibility scores for each entity according to Eq. 4 for each of the four query forms for six Pythia models of different sizes88870m, 410m, 1.4b, 2.8b, 6.9b, 12b (8-bit quantized) trained on the deduplicated Pile (Wolf et al., 2020; Dettmers et al., 2022; Biderman et al., 2023). Due to computational constraints, we approximate the model's answer distribution with the next token distribution over the model's vocabulary.999While it would be more precise to estimate the answer distribution by repeatedly sampling many model outputs, this is very computationally expensive. The detailed setup can be found in App. C. In the following sections, we describe further analyses and comparisons of these scores to validate the metrics and demonstrate how they can characterize model behavior.
4.2 Empirically Validating Our Metrics
4.2.1 Validating Persuasion Scores
Convergent Validity.
According to existing measurement modeling methods, observing a relationship between a new metric and existing ones would serve as additional evidence that the metric is meaningful. To this end, we explore the relationship between persuasion scores and whether the model outputted an answer consistent with the context. We expect that contexts with higher persuasion scores should also be more successful at convincing the model to answer in agreement with the context. To test this, we use the Pythia-6.9b-deduped model to generate a model answer for each prepended context to a query–entity pair from § 4.1. We map each answer to whether it agrees with the context, the original answer, or neither via simple string matching. We then run permutation tests () with the Benjamini-Hochberg (BH) correction (Benjamini and Hochberg, 1995) for the questions in each of the 122 query topics to test for whether contexts eliciting context-agreeing answers have higher persuasion scores than contexts which fail to change the answer from the original. Within a query topic, we run separate tests for open and closed queries, since the entropy of the answer distribution for closed queries tends to be much lower than the entropy for open queries.
We find that for of open queries, contexts that persuade the model to output the in-context answer have significantly higher persuasion scores than non-persuasive ones. Curiously, this behavior holds for only of closed queries. We discuss this surprising result more in § 6. We further show a summary of the mean persuasion scores for the different kinds of elicited answers in Fig. 2.
Construct Reliability.
To show construct reliability, we provide evidence that persuasion scores have low variance across different samples of contexts and different phrasings of the same query (i.e., query forms) for the Pythia-6.9b-deduped model. Since computing the persuasion score of a context requires marginalizing over all contexts (Eq. 3), we want to show that this metric is robust to different contexts sampled from the same distribution. The metric should also be robust to different phrasings of the same question, e.g., the persuasion score of The capital of Slovenia is Gotham City. should be similar when prepended to either Is Slovenia's capital Ljubljana? or Is Ljubljana the capital of Slovenia?.
[…]
4.2.2 Validating Susceptibility Scores
Convergent Validity.
We compare susceptibility scores to other metrics as further evidence for the meaningfulness of our metric. Since we are, to our knowledge, the first to conceptualize per-entity scores for susceptibility, we adapt the memorization ratio (MR) metric from Longpre et al. (2021) to apply on a per-entity basis, which is described in more detail in App. D. We compare each entity's susceptibility score and entity-specific MR for each query, using the same setup as in § 4.1 for the Pythia-6.9b-deduped model.
In Fig. 3, we note a decreasing upper bound relationship between susceptibility scores and MR for both open queries and closed queries. In order to interpret these results, it is important to note that these metrics differ in the nature of what they measure—susceptibility measures how the entropy of the output distribution is affected by context, whereas MR counts the actual outputs sampled from the model. Depending on the model's confidence without the prepended context, it is possible that both large and small changes in entropy (corresponding to large and small susceptibility scores) could indicate a change of output (low MR), while an unchanged output (high MR) indicates that the entropy likely remained similar (low susceptibility). Thus, we do not expect to see a straightforward correlation between the two metrics and instead expect a negative correlation between the upper bound of the susceptibility score at a given entity-specific MR. Both open and closed queries appear to exhibit this relationship (at least at the outliers), supporting the convergent validity of susceptibility. The open queries further show a decreasing pattern at each quartile for all bins except the lowest one (MR between 0 and 0.2). The different behavior for the lowest bin might be explained by a limitation of MR; especially for open queries, a model may fail to produce the original answer for most or all contexts if it does not know it from the start. In this case, MR cannot distinguish between the model behavior for these difficult queries, whereas susceptibility scores can provide more granular information about model behavior for such entities. This effect might be exacerbated by the majority of entities falling into the bin with the lowest MR for both open and closed queries. Additionally, the closed queries appear to have a mostly increasing pattern at each quartile across the bins. This could be explained by an artifact of the construction of the closed queries. Since the original answer of all closed queries is Yes, it is possible that the contexts increase the confidence in Yes due to some model artifact or token bias, which would explain higher susceptibility scores even for higher MR.
Construct Reliability.
Following the same setup used in § 4.2.1, we consider three seeds for sampling contexts and the same two query forms for both open and closed questions to test for reliability (low variance) for susceptibility scores across both axes of seeds/sampled contexts and query forms. Fig. 4 shows strong evidence for the reliability of susceptibility scores. Similarly to the persuasion scores, the variance for both open and closed queries is very low when taken across seeds and closed query forms; open query forms have higher variance because they have meaningfully different possible answers, as discussed in § 4.2.1.
4.3 What Makes a Context Persuasive?
To better characterize model behavior with persuasion scores, we explore several tests for qualities that might distinguish more persuasive contexts from less persuasive ones: relevance, assertiveness, and negation.
Permutation test results across models and comparisons
4.3.1 Relevance
Experiment Setup.
We use the setup described in § 4.1. For the relevance test, we consider a relevant context to be one that mentions the queried entity, and an irrelevant context as one that does not. We hypothesize that relevant contexts should be more persuasive than irrelevant ones. For each of the 122 queries, we find the mean persuasion score of the relevant contexts and irrelevant contexts, and use a permutation test to determine whether the mean persuasion score for relevant contexts (across entities) is higher than the mean persuasion score for irrelevant contexts, using a significance level of with the BH correction.
Results.
As seen in Fig. 5 (), across most model sizes and relations, relevant contexts are significantly more persuasive than irrelevant contexts. Specifically, depending on the model, 95–100% of open queries and 83–100% of closed queries showed a significant result. We also see a trend in which, as model size increases, so too does the degree to which relevant contexts are more persuasive (as measured by the mean effect size). We summarize the significance test results and effect sizes for all models and queries in App. E.
4.3.2 Assertiveness
Experiment Setup.
We use the same setup as in § 4.3.1 but instead explore our hypothesis that assertive contexts are more persuasive than the base ones by computing the mean persuasion score of each group and testing for a significant difference for each query.
Results.
Fig. 5 () shows that, consistently across model sizes, assertive contexts tend to be more persuasive than base contexts for closed queries but not for open queries. Moreover, assertive contexts are most persuasive for medium-sized models such as 1.4b and 2.8b. We hypothesize that smaller models may be less persuaded by assertive contexts because they may be worse at integrating context into its answer. Larger models may be less persuaded because of their stronger prior knowledge/lower susceptibility to context, whether assertive or not. We highlight the significance test results and effect sizes for all queries for the Pythia-6.9b-deduped model in Fig. 6 and other models in App. E.
4.3.3 Negation
Experiment Setup.
We use the same setup as in § 4.3.2 and explore whether negation contexts differ in persuasiveness from the base ones, using a two-tailed permutation test for each query.
Results.
The permutation tests suggest evidence for a significant difference in 88% of the closed queries and 74% of the open queries for the Pythia-6.9b-deduped model. However, there is not a consistent directional pattern; from Fig. 7, we see that negations are significantly more persuasive for some queries while significantly less persuasive for others. This pattern is similar for other models as seen in App. E. Fig. 5 () also shows that the smallest model is the most sensitive to being persuaded differently by negation vs base contexts. We hypothesize this may be due to potential spurious correlations/token biases between, for example, seeing the word not and the model's probability of outputting No.
4.3.4 Comparing Context Qualities
For all models on open queries, the relevance of a context has the greatest effect on persuasion score compared to assertiveness or negation (as measured by effect size). This is consistent with our intuition that the model should be more sensitive to whether the queried entity is mentioned in the context than to other context features. However, we can see in Fig. 5 that for medium-sized models on closed queries, the other context comparisons have stronger effect sizes (e.g., assertiveness and negation for the 2.8b model). This could be explained by potential spurious correlations/token biases associating, i.e., the word not with the model outputting No or the word definitely with outputting Yes.
4.4 What Makes an Entity Susceptible?
4.4.1 Familiar vs Unfamiliar Entities
We hypothesize that the model should have lower susceptibility scores for entities encountered during pretraining compared to unfamiliar, fake entities.
Experiment Setup.
We use the setup described in § 4.1 to compute susceptibility scores for known and unknown entities. We use a permutation test with and the BH correction to test for whether the known real entities are less susceptible than the unknown fake entities.101010We filter out fake entities that appear in the training data to better represent unknown entities in our analysis. The detailed setup can be found in App. C.
Results.
For the Pythia-6.9b-deduped model, we find that with queries (open questions) and queries (closed questions), familiar entities have significantly lower susceptibility scores than unfamiliar fake entities. We conjecture that for the remaining queries, the model may not have strong prior biases about the sampled entities for these queries. Indeed, further analysis (§ F.1) finds some evidence supporting the hypothesis that queries with smaller effect sizes/less significant -values feature less familiar entities, as we find a small correlation between effect size and entity frequency in the training set. Fig. 8 shows the distribution of susceptibility scores for real and fake entities for an example query with a particularly strong effect size.
Generally, as model size increases, so too does the significance and effect size of unfamiliar entities being more susceptible than familiar entities, as seen by the generally increasing blue lines () in all four plots in Fig. 5. This trend is consistent with our expectation that bigger models have stronger prior knowledge of entities and are therefore less susceptible to familiar entities. The smallest model (70m) does not have a significant difference in susceptibility between familiar and unfamiliar entities, which could indicate it is too small to have strong prior knowledge.
4.4.2 Degrees of Familiarity
Training Data Frequency.
Since language models are parameterized with knowledge from their training corpora, we hypothesize that the model is less susceptible for entities with which it is more familiar, i.e., more frequently occurring in the training data. We investigate this relationship between the Pythia models behavior and frequency statistics in the Pile dataset on which they were trained Gao et al. (2020). In an attempt to capture the model's familiarity with the relation between the entity and answer, we count the number of co-occurrences between the entity and its corresponding answer within a 50-word window. We compare the susceptibility score to this co-occurrence frequency and find a significant correlation (Spearman ) for the Pythia-6.9b-deduped model. We see in Fig. 9 that as the training data frequency increases, the susceptibility scores' upper bound decreases. This trend is shared across all model sizes (see § F.2).
Entity Degree in a Knowledge Graph.
Training data can be noisy, difficult, and expensive to search through for entity co-occurrences and more complex frequency statistics. Knowledge graphs offer a more precise alternative as they represent entities and relations extracted from common corpora like Wikipedia in a structured way. For example, within the pretraining data, it is very difficult to identify the number of different answers with which an entity will co-occur within the context of a specific relation. However, with a knowledge graph, we can easily identify the exact number of objects with which an entity might share a given relation. Thus, we explore the relationship between the relation-dependent degree of an entity in a knowledge graph and the susceptibility score. Similarly to the comparison against training data frequency, Fig. 10 also appears to show a decreasing upper bound between susceptibility scores and the degree for the Pythia-6.9b-deduped model; § F.2 further shows this trend is shared across all model sizes. We also provide the detailed methodology in § F.2. Such a trend suggests that unfamiliar entities have the potential to be highly susceptible, while very familiar entities tend to have lower susceptibility.
5 Applications
We examine how knowing susceptibility scores can be useful for analyzing model behavior in two different applications. Here, we highlight one key finding per application; however, we emphasize that future work can conduct a more detailed analysis of these and other applications.
Social Sciences Measurement.
Social scientists use large language models (LLMs) for annotating data and descriptive data analysis, yet such use may inadvertently incorporate entity biases and skew model behavior (Ziems et al., 2024; Gilardi et al., 2023; Zhang et al., 2023; O'Hagan and Schein, 2023). To better contextualize model annotations for a case study on friend-enemy stance detection (Choi et al., 2016; Stoehr et al., 2023), we aim to understand how susceptibility scores may differ between friend and enemy-based entity pairs for the query The relationship between {entity1} and {entity2} is, e.g., are famous friend-based relationships more susceptible than enemy-based relationships? From Fig. 11, we see that with the Pythia-6.9b-deduped model for two specific query forms, enemy duos are less susceptible than friend duos, which can inform social scientists that a model's annotation for friend pairs may be more easily influenced by the context than enemy pairs. We provide more details in App. G.
Exploring Gender Bias.
Since higher susceptibility scores indicate weaker induced biases for entities, we conjecture that this can relate to being underrepresented in the training data. Based on this, we consider how the susceptibility score can be used to study gender bias in LLMs. Using GPT-4, we collect stereotypically-biased contexts, gendered names, and neutral queries and run several experiments to identify gender discrepancies in susceptibility scores (see App. G for full details). We highlight a result where masculine names have a higher susceptibility than feminine names when swapping the genders in the stereotypical contexts, as seen in Fig. 12. This could indicate that the model is more surprised to see contexts claiming men follow feminine stereotypes, and therefore could suggest less representation of feminine stereotypes in the training data than masculine ones.
6 Discussion and Conclusion
We have made the case that the persuasion score and susceptibility score are both valid and reliable in measuring their respective constructs, following a well-established measurement modeling framework. Throughout our experiments, we find a common theme in the results: there is a strong, negative, possibly linear relationship between the upper bound of the susceptibility score and (a) the entity's memorization ratio (Fig. 3), (b) the log-co-occurrence frequency in the training data (Fig. 9), and (c) the log-relation-dependent degree in a knowledge graph (Fig. 10). That is, as each of those three values increases, we see a clear pattern indicating that the highest susceptibility scores tend to decrease. This is consistent with our hypothesis that the induced bias of an entity increases for a model as the model's expected familiarity with the entity increases. Furthermore, we find a difference in behavior between scores for open and closed queries; while in many experiments, we see similar patterns between the two, susceptibility and persuasion appear to have a stronger relationship with memorization ratio for open queries, while assertive contexts appear to be significantly more persuasive than base contexts primarily for closed queries. This difference is a surprising phenomenon that warrants future study; we hypothesize closed queries may behave this way due to common token biases (with its output space of Yes and No) and/or the influence of mentioning both an entity and an answer in the query. Finally, while we applied these metrics to analyze model behavior in two case studies, in future work we aim to apply them to unearth new perspectives in other context and prior knowledge-dependent problems such as retrieval-augmented generation, model editing and control, and few-shot learning.
Limitations
We face some technical limitations in executing the empirical aspects of this work. First, while § 3 defines the output space of as the set of all possible outputs , in practice, it is computationally expensive to estimate that probability distribution. Instead, we look only at the model's probability distribution of the next token, which could be a noisy signal, especially in cases where the answer suggested by a context and the answer suggested from prior knowledge share the same first token. Second, it is difficult to go through the whole Pile to count answer-entity co-occurrences without noise. Third, the scores depend on the sampled contexts, which may not be representative of all applications.
Ethics Statement
As LLM capabilities grow more advanced and their usage proliferates throughout the real world, we acknowledge that their development can exacerbate risks to people, especially those historically underrepresented or misrepresented to these models. Our work aims to make model behavior more transparent by providing a new tool to analyze the interaction between context and prior knowledge in LMs, which is especially important as people interact with them in chat, question-answering, and other prompt-based settings. We foresee no particular ethical concerns and hope this paper contributes to developing tools that can identify and mitigate ethical concerns in the future.
Acknowledgements
We thank Alex Warstadt, Tiago Pimentel, Anej Svete, Alexandra Butoi, Benjamin Dayan, and Leo Du for helpful comments, discussion, and feedback. Niklas Stoehr acknowledges funding through the Swiss Data Science Center (SDSC) Fellowship. Vésteinn Snæbjarnarson is supported by the Pioneer Centre for AI, DNRF grant number P1.
Appendix A A Primer on Half-pointwise Mutual Information
Half-pointwise mutual information (HPMI) is a non-standard concept. There is, for example, no mention of it in standard references on information theory (Cover and Thomas, 2006). In this brief primer, we give various properties of HPMI and show how it relates to other concepts in information theory. Given random variable over the discrete space , random variable over the discrete space , and , the half-PMI of is defined as:
such that
We now state and prove three Propositions about HPMI.
Proposition 1.
Given random variable over the discrete space , random variable over the discrete space , and , then:
Proof.
Proposition 2.
Given random variable over the discrete space , random variable over the discrete space , and , then:
where is the cross-entropy between and .
Proof.
Notably, . Furthermore, while the decomposition of mutual information into a difference in entropies is symmetric, i.e., , this shows that half-pointwise mutual information is not, i.e., .
Proposition 3.
Given random variable over the discrete space , random variable over the discrete space , and , then:
Proof.
Half-PMI is equivalent to by definition. ∎
Corollary 1.
Half-PMI is nonnegative.
Proof.
Since half-PMI is equivalent to (Proposition 3) and KL-divergence is nonnegative, half-PMI must be non-negative. ∎
Appendix B An Entity-Independent Persuasion Score
In addition to the persuasion score, which is entity-dependent, we also consider an entity-independent extension. We assign an entity-independent persuasion score to a context which represents how convincing or persuasive a context is at altering a model's answer distribution to a query, regardless of which entity parameterizes the query. One might be interested in an entity-independent persuasion score for contexts like Only give wrong answers to questions., which we might expect to have an effect on all queries regardless of the entity. Another use case is in comparing the persuasiveness of context templates, such as {entity1} loves {entity2} and {entity1} really really really loves {entity2} for the query What's the relationship between {entity1} and {entity2}?. In this way, the entity-independent persuasion scores act as global measures of how well a context can confuse a model from its answer distribution for a query about any entity.
Analogous to our definition of the susceptibility score, we define the entity-independent persuasion score of a context as how much the log probability distribution of possible answers changes, averaged across all possible entities and answers. More precisely, we define our entity-independent persuasion score as:
which is the half-conditional PMI. Further marginalizing out the context, we arrive at the entity-independent susceptibility score for a query:
The entity-independent persuasion score differs from the persuasion score by additionally marginalizing over the entities. As the half-PMI is conditioned on the entity random variable, this tells us, when we already know the entity, for a given context, how much more confident can we be in the answer. In some sense, then, this can be interpreted as the average persuasiveness of a context across all entities for the query.
Appendix C Detailed Experimental Setup
We extract 122 relations from the YAGO knowledge graph Suchanek et al. (2007), such as alumniOf, capital, and highestPoint. For each relation, we do the following:
We randomly sample real entities (and corresponding answers) from YAGO and use GPT-4 111111gpt-4-1106-preview, January 2024 (OpenAI, 2023) to generate fake entities with the same entity class as the real ones121212 Entity classes: CreativeWork, Event, Intangible, Organization, Person, Place, Product, Taxon, and FictionalEntity..131313Real example: Adele. Fake example: Udo König.
[…]
We construct context templates of 3 types: base (e.g., The capital of {entity} is {answer}.), assertive (e.g., The capital of {entity} is definitely {answer}.), and negation (e.g., The capital of {entity} is not {answer}.). We parameterize these context templates with both real and fake entities (and answers, if applicable). From this, we randomly sample contexts, subject to the constraint that each entity is directly mentioned in 6 contexts total (that is, in 2 assertive contexts, 2 base contexts, and 2 negation contexts).
We compute the persuasion scores for the real and fake entities according to Eq. 2.
We compute the susceptibility score for the real and fake entities according to Eq. 4. We approximate with a uniform distribution over the set of sampled contexts and with the model's next token probabilities.
For the various group comparisons (e.g., relevant vs irrelevant contexts, familiar vs unfamiliar entities, etc.), we use a permutation test over the t-statistic (, with the BH correction) to test our null hypothesis for each comparison.
Appendix D Entity-Specific Memorization Ratio
The memorization ratio (MR), as used by Longpre et al. (2021), is defined as follows: Given a set of ( query, context)-pairs with knowledge conflicts (i.e., the answer in the context disagrees with the original answer), , where is the number of queries for which the model returned the original answer and is the number of queries for which the model returned the substitute answer presented in the context. Then, the entity-specific MR follows this definition under the additional constraint that the query/entity is fixed and we vary only the contexts; the resulting number tells us, for a given query about an entity, the fraction of contexts for which the model returned the original answer instead of the substitute one.
Appendix E Persuasion Scores: In-Depth Results
E.1 Relevant vs Irrelevant Context Persuasion Scores Across Models
Our null hypothesis that the mean persuasion score of relevant contexts is not greater than that of irrelevant contexts. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 13.
E.2 Assertive vs Base Context Persuasion Scores Across Models
Our null hypothesis that the mean persuasion score of assertive contexts is not greater than that of base contexts. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 14.
E.3 Negation vs Base Context Persuasion Scores Across Models
Our null hypothesis that the mean persuasion score of negation contexts is not equal to that of base contexts. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 15.
Appendix F Susceptibility Scores: In-Depth Results
F.1 Unfamiliar vs Familiar Entity Susceptibility Scores Across Models
Our null hypothesis that the mean susceptibility score of unfamiliar entities is not greater than that of familiar entities. We summarize the test results (effect size and p-values) for all queries for all models in Fig. 16. From this figure, we can see the trend of how effect size and percentage of significant queries generally increases with model size, and notably the smallest model has no significant results for any query. However, even the larger models do not exhibit significant differences in scores between unfamiliar and familiar entities for all queries. To investigate the spread further, we plot the p-values and effect size against the entity frequencies in the Pile. The results are presented in Fig. 17. There is a significant trend for the open queries against the frequency (spearman, is , ), showing that real entities tend to be less susceptible the more frequently they appear in the training data. The trend is not significant for the closed set.
F.2 Training Data and Susceptibility Scores
Since language models are parameterized with knowledge from their training corpora, we examine whether we can identify correlations between patterns in entity susceptibility scores and their prevalence in the training data.
Since the Pythia models are all trained on the Pile dataset Gao et al. (2020) in a single pass, we choose to compare the susceptibility scores from the Pythia models to various frequency and co-occurrence statistics in the Pile.141414We use the same deduplicated 825GiB version of the Pile that the Pythia-6.9b model was trained on. https://huggingface.co/datasets/EleutherAI/the_pile_deduplicated.
Experiment Setup.
Our goal is to understand how the susceptibility score relates to the frequencies of entities and their co-occurrences in the training data. For this, we use all of the entities and answers selected as described in § 4.4 and locate them in the Pile. We only perform rudimentary tokenization of each document by removing punctuation and splitting at white spaces, but find this suffices to locate the exact terms. As a sanity check, we annotate named entities in 30k documents and cross-reference the list of entities. If a supposed entity has a fairly high frequency (>50) and is most often (>75%) not labeled as an entity, we exclude it from the calculations. This removes 200 entities that are high-frequency non-entity words in English and lowers the co-occurrences by 25%. Finally, we calculate the token distance for each entity-answer pair for every document in 1/3 of the Pile.
Results.
We compare the co-occurrences of the entity-answer pairs to the averaged susceptibility scores over all queries that apply to the entity , . Our results show a stark difference in behavior between the open and closed questions (§ 4.4). The open questions not only have higher susceptibility scores, but the training corpus frequency of the mentioned entities influences them more. This is to be expected as there are far more probable candidates for open questions than for closed yes–no-style questions. We also find a significant negative correlation (Spearman -0.23, ) between frequency and susceptibility scores for the pythia-6.9b-deduped model, indicating that the language model is less susceptible to context interference for entity-answer pairs that are more frequently found in the training corpus. See Fig. 9 and Fig. 18. Finally, we also notice a big difference in the rank correlation depending on the query type. Some query types are more susceptible to context for the given entities than others. An overview of this is given in the next section.
Susceptibility and Frequency Analysis
We are interested in seeing how different predicate relations of the queries from the knowledge graph have different susceptibility scores. We evaluate the relation between susceptibility scores and co-occurrence frequencies per entity-answer predicate relation. This reveals trends in what types of relations are more susceptible to context than others. A lower correlation indicates a stronger drop in susceptibility as the entities become more frequent in the pretraining data.
F.3 Entity Degrees and Susceptibility Scores
Since knowledge graphs are structured conceptual maps relating entities, we further seek to identify whether we can identify any correlations between statistics of the YAGO knowledge graph and an entity's susceptibility scores.
Validating against a knowledge graph can be advantageous over validating against the actual training data for a number of reasons, including that for most models, the actual training data is inaccessible for research, and the scale of the training data can make it prohibitively expensive to trawl through efficiently and precisely. For example, with a knowledge graph, we can identify the exact number of objects an entity might share an alumniOf relation with, while within the pretraining data, it is very difficult to identify the number of different answers an entity will co-occur within the context of the specific alumniOf relation.
Experiment Setup.
Our goal is to understand how the susceptibility score relates to the degree of entities in the YAGO knowledge graph for specific queries. For this, we extract the number of incoming and outgoing edges from an entity along a relation (or query) as follows: . We plot against the susceptibility score for all entities and queries.
Results.
From Fig. 19, we see a decreasing upper bound relationship between susceptibility scores and the YAGO degree for both open and closed queries for all model sizes. This could be explained as follows: consistent with our original hypothesis, very familiar entities to a model have low susceptibility, while less familiar entities can have a wider range of susceptibility. The potential for unfamiliar entities to be susceptible is much higher than that for very familiar entities, although unfamiliar entities can also be less susceptible. Further investigation into traits that characterize the susceptibility of familiar vs less familiar entities is needed.
Appendix G Applications
G.1 Social Sciences Measurement
Motivation.
Large language models (LLMs) are actively used today in empirical social sciences for annotating data and descriptive data analysis (e.g., classifying tweets with sentiment and ideology scores (Ziems et al., 2024; Gilardi et al., 2023)). However, Zhang et al. (2023) warn that LLMs applied to sentiment classification ``may inadvertently adopt human biases'' and demonstrate that the prompt design, i.e., the context persuasiveness, can significantly influence the outcome. O'Hagan and Schein (2023) demonstrate that LLMs exhibit biases about different entities when measuring political ideology, based on their prior knowledge. Finally, Stoehr et al. (2024) while use LLMs to measure the stance of product reviews, their setting does not disentangle the effects of prior knowledge and context, thus leaving ambiguous the question of whether the measurement is more because of the LLM's prior bias about the product or the review's actual content.
Experiment Setup.
We consider a manually constructed dataset (Tab. 1) of well-known entity-pairs which are either friends (e.g., Harry Potter and Ron Weasley) or enemies (e.g., David and Goliath) and contexts relating the pair (e.g., Harry loves Ron). We aim to understand how susceptibility scores may differ between the two kinds of relationships for the query What's the relationship between {entity1} and {entity2}?, e.g., are famous friend-based relationships more susceptible than enemy-based relationships? We compute the susceptibility scores for these entity-pairs using simple template-generated contexts such as {entity1} loves {entity2} and {entity1} hates {entity2} and then analyze the results.
Results.
From Fig. 11, we can see that for the open query The relationship between {} and {} is, there is a clear difference in susceptibility scores for friend and enemy entity-pairs. We leave to future work investigating the exact nature of why some entity-pairs may have lower susceptibility scores than others.
G.2 Exploring Gender Bias
Motivation.
The susceptibility score tells us how the entropy changes for a given entity as we vary the prepended contexts. We posit that stronger induced biases for entities lead to lower susceptibility scores than for those less represented in the training data. Using this information, we consider how the susceptibility score can be used to study gender bias in LLMs.
Experiment Setup.
We collect sentences containing 20 highly stereotypical sentences about men and women, respectively, to use as contexts. For instance, Men are not good at multitasking. We also collect 20 names for each gender and 20 neutral queries that we think may be affected by the contexts, such as {entity} spends most of their time. Finally, we also invert the queries by flipping the gendered words, giving us non-stereotypical contexts. We then calculate susceptibility scores for each query and gender combination over the four different context collections. All scores are calculated using the Pythia-6.9b-deduplicated model. All data was generated with GPT-4. The contexts, names, and queries can be found below.
Results.
Results are shown in Fig. 12. For all contexts, there is a significant difference between the genders (). We make several observations: (i) When prompted with the masculine stereotypes, we get the lowest susceptibility for the male entities, indicating the model has a bias for the stereotypical male contexts. (ii) By swapping the genders in the stereotypical female contexts (M*), we get the highest susceptibility for the male entities. Somewhat surprisingly, we also find that this gives the lowest susceptibility scores for the female names. (iii) The susceptibility scores for the female names over the stereotypical female contexts (F) are similar to the swapped ones (F*), hinting that the model does not have as strong a bias for the feminine stereotypes as in the masculine case in (i). Combined with our prior findings on the relation between susceptibility and the saliency of entities, this hints at masculine stereotypes being more prominently represented in the training data (as well as in the model's induced biases). At the same time, we don't see evidence of exaggerated feminine stereotypes being favoured via lower susceptibility scores.
G.2.1 Gender Stereotypes Data
Here, we list the data used to run the gender bias use-case. All data was generated with the help of the GPT-4 LLM.
The female names are Mary, Jennifer, Linda, Patricia, Elizabeth, Susan, Jessica, Sarah, Karen, Nancy, Margaret, Lisa, Betty, Dorothy, Sandra, Ashley, Kimberly, Donna, Emily, Michelle. The male names are James, John, Robert, Michael, William, David, Richard, Joseph, Thomas, Charles, Christopher, Daniel, Matthew, George, Donald, Anthony, Paul, Mark, Andrew, Edward.
The stereotypical feminine contexts are: Women are not as strong as men., All women are overly emotional and sensitive., Women are not good at math and sciences., All women are naturally good at cooking and cleaning., Women talk more than men., Women are generally not capable of doing manual labor., All women love shopping and fashion., Women don't understand technology as well as men do., Women are indecisive and change their mind too often., Women cannot drive as well as men., Women are not as ambitious as men when it comes to career., Women overthink things more than men., Women are not as good at sports as men., Women are more gossipy than men., All women are afraid of bugs and spiders., Women are not good at directions and get lost easily., Women take longer to get ready than men., Women always need to ask for others for their opinion., All women are obsessed with their weight and dieting., All women are crazy about babies and want to have kids.
The stereotypical masculine contexts are Men always bring the association of strength and aggression., All men are inherently messy and cannot keep their spaces clean., Men are not good at multitasking., All men are naturally good at sports., Men cannot cook as well as women., Men do not take care of their health as much as women do., I believe men are not as talkative as women., All men love video games., Men are not good listeners., Men are always competitive., Men are not as emotional as women., Men are not able to express their feelings as well as women do., Men only think about one thing., Men look at problems from a logical perspective, not considering emotions., Men are typically less organized than women., Men tend to make decisions without considering others' feelings., All men only care about sports and cars., Men don't like to ask for help., Men don't pay as much attention to appearance as women do., All men can't resist a flirt.
The queries used for both sets of entities are: I think {entity} works as a, {entity} is good at, I met {entity} yesterday, they are, {entity} usually spends their free time, {entity}'s favorite hobby is likely, {entity} probably lives in a, When it comes to {entity}'s fashion style, they prefer, {entity}'s educational background might be in, {entity} often enjoys eating, If {entity} was a colour, they would be, The one thing that {entity} cannot resist is, {entity} probably drives a, One thing that {entity} is really bad at is, {entity}'s personality can be described as, {entity} spends most of their time, {entity} probably likes to watch, During weekends, {entity} might be, {entity}'s favorite type of music is probably, I imagine {entity} has a pet, it is a, {entity}'s favorite type of movie is probably.
| Entity 1 | Entity 2 | Label | Type |
| Sherlock Holmes | James Moriarty | enemy | movie |
| Obi Wan Kenobi | Darth Maul | enemy | movie |
| Harry Potter | Lord Voldemort | enemy | movie |
| Harry Potter | Draco Malfoy | enemy | movie |
| Spiderman | Norman Osborne | enemy | movie |
| Super Mario | Bowser | enemy | movie |
| Gandalf | Saruman | enemy | movie |
| Bilbo Baggins | Sauron | enemy | movie |
| Superman | Lex Luthor | enemy | movie |
| James Bond | Ernst Stavro Blofeld | enemy | movie |
| Optimus Prime | Megatron | enemy | movie |
| Boston Red Sox | New York Yankees | enemy | sports |
| Green Bay Packers | Chicago Bears | enemy | sports |
| Borussia Dortmund | FC Bayern Munich | enemy | sports |
| Real Madrid | FC Barcelona | enemy | sports |
| Joe Frazier | Muhammad Ali | enemy | sports |
| AC Milan | Inter Milan | enemy | sports |
| Torries | Labor Party | enemy | politics |
| Democrats | Republicans | enemy | politics |
| USA | Al-Qaeda | enemy | politics |
| Donald Trump | Hillary Clinton | enemy | politics |
| Donald Trump | Joe Biden | enemy | politics |
| Kuomintang | Chinese Communist Party | enemy | history |
| Winston Churchill | Adolf Hitler | enemy | history |
| Harry Trumann | Nikita Khrushchev | enemy | history |
| George Bush | Saddam Hussein | enemy | history |
| David | Goliath | enemy | history |
| Greece | Troy | enemy | history |
| Gauls | Rome | enemy | history |
| USA | Soviet Union | enemy | history |
| Nazi Germany | Allied Forces | enemy | history |
| Cain | Abel | enemy | history |
| Coca Cola | Pepsi | enemy | business |
| Ford | General Motors | enemy | business |
| Thomas Edison | Nikola Tesla | enemy | business |
| Steve Jobs | Bill Gates | enemy | business |
| Airbus | Boeing | enemy | business |
| McDonalds | Burger King | enemy | business |
| Visa | Mastercard | enemy | business |
| Netscape | Microsoft Internet Explorer | enemy | business |
| UPS | Fedex | enemy | business |
| Canon | Nixon | enemy | business |
| Sony | Nintendo | enemy | business |
| Sheriff of Nottingham | Robin Hood | enemy | history |
| Moby Dick | Captain Ahab | enemy | movie |
| Tom | Jerry | enemy | movie |
| Peter Pan | Captain Hook | enemy | movie |
| Jack Sparrow | Hector Barbossa | enemy | movie |
| Harry Potter | Ronald Weasley | friend | movie |
| John Lennon | Paul McCartney | friend | history |
| Amelia Earhart | Eleanor Roosevelt | friend | history |
| Georges Braque | Pablo Picasso | friend | history |
| Bilbo Baggins | Gandalf | friend | movie |
| Harry Potter | Hermione Granger | friend | movie |
| Frodo Baggins | Samwise Gamgee | friend | movie |
| Han Solo | Chewbacca | friend | movie |
| C.S. Lewis | J.R.R. Tolkien | friend | history |
| Alexander the Great | Hephaestion | friend | history |
| Mark Twain | Nikola Tesla | friend | history |
| John Adams | Thomas Jefferson | friend | history |
| Bill Gates | Warren Buffett | friend | business |
| Vincent van Gogh | Paul Gauguin | friend | history |
| Albert Einstein | Niels Bohr | friend | history |
| Woody | Buzz Lightyear | friend | movie |
| Shrek | Donkey | friend | movie |
| Bal Gangadhar Tilak | Mohammed Ali Jinnah | friend | history |
| Marc Twain | Hellen Keller | friend | history |
| Thomas Edison | Henry Ford | friend | history |
| Bill Gates | Paul Allen | friend | business |
| Larry Page | Sergei Brin | friend | business |
| Mike Wazowski | James P. Sullivan | friend | movie |
| Sherlock Holmes | John Watson | friend | movie |
| Harry Potter | Albus Dumbledore | friend | movie |